### v3

@@ 1 Introduction
    Extension of variable smoothing algorithms for Problem 1.2 with asymptotic convergence analysis. For finding a stationary point of Problem 1.1 via Problem 1.2, we extend a variable smoothing algorithm [23] (see Algorithm 1 in Section 4.2), where the proposed algorithm is designed as a single-loop algorithm. For the proposed algorithm, Theorem 4.9 (c) presents an asymptotic convergence analysis in terms of a stationary point, with the aid of the gradient consistency property of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$. Moreover, Theorem 4.9 (d) presents an iteration complexity of $\mathcal{O}(\epsilon^{-3})$ to obtain an $\epsilon$-approximate stationary point (see also Remark 4.10). Our asymptotic convergence analysis can also apply to the existing variable smoothing algorithm [23] for the unconstrained case because [23] is recovered by the proposed algorithm with a linear operator $\mathfrak{S}$ and $\bm{\varphi}\coloneqq\mathrm{Id}$. Finally, numerical experiments, in Section 5, demonstrate the efficacy of the proposed algorithm in two scenarios of the sparse spectral clustering and the sparse principal component analysis.
@@
For a differentiable function $J:\mathcal{X}\to\mathbb{R}$, $\nabla J(\bm{x})\in\mathcal{X}$ is the gradient of $J$ at $\bm{x}\in\mathcal{X}$ if $\mathrm{D}J(\bm{x})[\bm{v}]={\left\langle\nabla J(\bm{x}),\bm{v}\right\rangle}$ for all $\bm{v}\in\mathcal{X}$. Note that for a differentiable mapping $\mathcal{F}:\mathcal{Y}\to\mathcal{X}$, we have the expression [^2] It can be checked by $(\forall\bm{v}\in\mathcal{Y})\ {\left\langle\nabla(J\circ\mathcal{F})(\bm{x}),\bm{v}\right\rangle}_{\mathcal{Y}}=\mathrm{D}(J\circ\mathcal{F})(\bm{x})[\bm{v}]=\mathrm{D}J(\mathcal{F}(\bm{x}))[\mathrm{D}\mathcal{F}(\bm{x})[\bm{v}]]={\left\langle\nabla J(\mathcal{F}(\bm{x})),\mathrm{D}\mathcal{F}(\bm{x})[\bm{v}]\right\rangle}_{\mathcal{X}}={\left\langle(\mathrm{D}\mathcal{F}(\bm{x}))^{*}[\nabla J(\mathcal{F}(\bm{x}))],\bm{v}\right\rangle}_{\mathcal{Y}}$. $\nabla(J\circ\mathcal{F})(\bm{y})=(\mathrm{D}\mathcal{F}(\bm{y}))^{*}[\nabla J(\mathcal{F}(\bm{y}))]\ (\bm{y}\in\mathcal{Y})$. For a function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, $\mathrm{dom}(J)\coloneqq\{\bm{x}\in\mathcal{X}\mid J(\bm{x})<+\infty\}\subset\mathcal{X}$ and $\mathrm{epi}(J):=\{(\bm{x},a)\in\mathcal{X}\times\mathbb{R}\mid J(\bm{x})\leq a\}\subset\mathcal{X}\times\mathbb{R}$ denote respectively the domain and the epigraph of $J$. Moreover, $J$ is said to be (i) proper if $\mathrm{dom}(J)\neq\emptyset$; (ii) lower semicontinuous if $\mathrm{epi}(J)$ is closed in $\mathcal{X}\times\mathbb{R}$; (iii) locally Lipschitz continuous at $\widebar{\bm{x}}\in\mathcal{X}$ if there exist an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ and $L>0$ such that $\lvert J(\bm{x}_{1})-J(\bm{x}_{2})\rvert\leq L\left\lVert\bm{x}_{1}-\bm{x}_{2}\right\rVert\ (\bm{x}_{1},\bm{x}_{2}\in\mathcal{N}_{\widebar{\bm{x}}})$. For a function $J:\mathcal{X}\to\mathbb{R}\cup\{-\infty,+\infty\}$, the limit inferior [57, Definition 1.5] and the limit superior [57, p.13] of $J$ at $\widebar{\bm{x}}\in\mathcal{X}$ are defined respectively by
@@
For a set sequence $(E_{n})_{n=1}^{\infty}$ with subsets $E_{n}\subset\mathcal{X}$, the outer limit [57, Definition 4.1] of $(E_{n})_{n=1}^{\infty}$ is defined by
@@
where $\aleph_{0}$ stands for the cardinality of $\mathbb{N}$. With the outer limit of a set sequence, the outer limit [57, p.152] of a set-valued mapping $S:\mathcal{X}\rightrightarrows\mathcal{X}$ at $\widebar{\bm{x}}\in\mathcal{X}$ is defined as
@@ 2.1 Subdifferential and Normal Cone
where $\partial_{\rm F}J(\widebar{\bm{x}})\coloneqq\left\{\bm{v}\in\mathcal{X}\mid\liminf_{\mathcal{X}\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{J(\bm{x})-J(\widebar{\bm{x}})-{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\geq 0\right\}$ is called the Fréchet (or regular) subdifferential of $J$ at $\widebar{\bm{x}}\in\mathrm{dom}(J)$, and $\partial_{\rm L}J(\widebar{\bm{x}})$ and $\partial_{\rm F}J(\widebar{\bm{x}})$ at $\widebar{\bm{x}}\notin\mathrm{dom}(J)$ are understood as $\emptyset$ (Note: for a convex function $J$, $\partial_{\rm L}J$ and $\partial_{\rm F}J$ are identical to the standard convex subdifferential of $J$ [57, Prop. 8.12]).
By using the subdifferentials, the general normal cone $N_{C}(\widebar{\bm{x}})$ and the regular normal cone $\widehat{N}_{C}(\widebar{\bm{x}})$ to a nonempty subset $C\subset\mathcal{X}$ at $\widebar{\bm{x}}\in C$ are defined (see [57, Def. 6.3, Prop. 6.5, and Ex. 8.14]) respectively by
@@
With notions of normal cone operators, we define regularities of a subset and functions as follows. Such regularities ensure the chain and sum rules for the limiting subdifferential as shown in Lemma 2.4 below (see also Fact A.4 in Appendix A).
@@
    (Clarke regularity [57, Def. 6.4]) A nonempty set $C\subset\mathcal{X}$ is said to be Clarke regular at $\widebar{\bm{x}}\in C$ if (i) $N_{C}(\widebar{\bm{x}})=\widehat{N}_{C}(\widebar{\bm{x}})$ (Note: the inclusion $N_{C}(\widebar{\bm{x}})\supset\widehat{N}_{C}(\widebar{\bm{x}})$ always holds), and (ii) $C\cap\overline{\mathcal{N}}_{\widebar{\bm{x}}}$ is closed for some closed neighborhood [^3] In other words, $\overline{\mathcal{N}}_{\widebar{\bm{x}}}$ is a closed set that contains $\widebar{\bm{x}}$ as an interior point of $\overline{\mathcal{N}}_{\widebar{\bm{x}}}$. $\overline{\mathcal{N}}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$. Moreover, $C$ is said to be Clarke regular if it is Clarke regular at every $\widebar{\bm{x}}\in C$.
@@
    (Subdifferential regularity [57, Def. 7.25]) A proper lower semicontinuous function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ is said to be subdifferentially regular at $\widebar{\bm{x}}\in\mathrm{dom}(J)$ if $\mathrm{epi}(J)\subset\mathcal{X}\times\mathbb{R}$ is Clarke regular at $(\widebar{\bm{x}},J(\widebar{\bm{x}}))$. Moreover, $J$ is said to be subdifferentially regular if it is subdifferentially regular at every $\widebar{\bm{x}}\in\mathrm{dom}(J)$. We note that all functions in this paper are shown to be subdifferentially regular (see, e.g., Lemma 2.4 below and also Example A.2).
A typical Clarke regular set is a closed convex set [57, Thm. 6.9]. Another important example of Clarke regular sets is an embedded submanifold of the Euclidean space $\mathcal{X}$ [57, Exm. 6.8] as follows.
Example 2.3 (Embedded submanifold of $\mathcal{X}$ [57, Exm. 6.8]).
Let $(\emptyset\neq)C\subset\mathcal{X}$ be a $d(\leq\dim(\mathcal{X}))$-dimensional embedded submanifold of $\mathcal{X}$, i.e., for each $\widebar{\bm{x}}\in C$, there exist (i) an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ in $\mathcal{X}$ and (ii) a continuously differentiable mapping $H_{\widebar{\bm{x}}}:\mathcal{N}_{\widebar{\bm{x}}}\to\mathbb{R}^{\dim(\mathcal{X})-d}$, such that $C\cap\mathcal{N}_{\widebar{\bm{x}}}=H_{\widebar{\bm{x}}}^{-1}(\bm{0})$ and $\mathrm{Rank}(\mathrm{D}H_{\widebar{\bm{x}}}(\widebar{\bm{x}}))=\dim(\mathcal{X})-d$ (Note: by Whitney’s embedding theorem (see, e.g., [62, Thm. 6.15]), every $d$-dimensional smooth manifold can be seen as an embedded submanifold in $\mathbb{R}^{2d+1}$). Then, by [57, Def. 6.4 and Exm. 6.8], the embedded submanifold $C$ is Clarke regular, and we have the following expression:
@@
where the orthogonal complement subspace $T_{C}(\widebar{\bm{x}})$ of $N_{C}(\widebar{\bm{x}})$ is called the tangent cone to $C$ at $\widebar{\bm{x}}$ (Note: $N_{C}(\widebar{\bm{x}})$ and $T_{C}(\widebar{\bm{x}})$ are also called respectively the normal space and the tangent space to the manifold $C$ at $\widebar{\bm{x}}\in C$ [32, Thm. 3.15, Def. 3.16, and p.107]).
Lemma 2.4 (Subdifferential calculus for Problems 1.1 and 1.2).
Consider $F=h+g\circ\mathfrak{S}$, $C\subset\mathcal{X}$ and $\bm{\varphi}$ in Problems 1.1 and 1.2. Then, $F$, $F+\iota_{C}$ and $F\circ\bm{\varphi}$ are subdifferentially regular, and we have the following expressions:
@@
Fact 2.5 (Normal cone to image sets [57, Thm. 6.43]).
@@
The first inclusion is verified by [57, Thm. 6.43]. The equality follows by the relation $(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\bm{v}]\in\{\bm{0}\}\Leftrightarrow\bm{v}\in\mathrm{Ker}\left((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}\right)\ (\widebar{\bm{y}}\in\bm{\varphi}^{-1}(\widebar{\bm{x}}))$. ∎
@@ 2.2 Proximity Operator and Moreau Envelope
The proximity operator and the Moreau envelope have been serving as key ingredients for minimization of nonsmooth convex functions (see, e.g., [6, 7, 8]). Even for weakly convex functions, they are useful in design of nonsmooth optimization algorithms. We note that closed-form expressions of the proximity operators for commonly-used weakly convex functions have been found, e.g., $\ell_{1}$-norm [6], Minimax Concave Penalty (MCP) [16], Smoothly Clipped Absolute Deviation (SCAD) [63], log-sum penalty [64], piece-wise exponential function [65], and arctangent penalty [66].
Definition 2.6 (Proximity operator and Moreau envelope [57, Def. 1.22]).
@@
    For any $\widebar{\bm{z}}\in\mathcal{Z}$ and $(\bm{z}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ satisfying $\lim_{n\to\infty}\bm{z}_{n}=\widebar{\bm{z}}$, we have $\lim_{n\to\infty}\mathrm{prox}_{\mu_{n}g}(\bm{z}_{n})=P_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}(\widebar{\bm{z}})$ if $(\mu_{n})_{n=1}^{\infty}\searrow 0$ with $\mu_{n}\in(0,\eta^{-1})$ [67, Lemma 2.3], where $\mathrm{dom}(\partial_{\rm L}g)\coloneqq\{\bm{z}\in\mathcal{Z}\mid\partial_{\rm L}g(\bm{z})\neq\emptyset\}\subset\mathcal{Z}$, $\overline{(\cdot)}$ denotes the closure of a given set, and $P_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}=\mathrm{prox}_{\iota_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}}$ denotes the projection mapping onto the closed convex set $\overline{\mathrm{dom}(\partial_{\rm L}g)}\subset\mathcal{Z}$.
@@
    For any $\mu\in(0,\eta^{-1})$, (i) $^{\mu}g:\mathcal{Z}\to\mathbb{R}$ is continuously differentiable with $\nabla{}^{\mu}g=\frac{\mathrm{Id}-\mathrm{prox}_{\mu g}}{\mu}$, and (ii) $\nabla{}^{\mu}g$ is Lipschitz continuous with a Lipschitz constant $\max\left\{\mu^{-1},\frac{\eta}{1-\eta\mu}\right\}$ [68, Cor. 3.4].
@@
    $(\widebar{\bm{z}}\in\mathcal{Z},\mu\in(0,\eta^{-1}))\quad\nabla{}^{\mu}g(\widebar{\bm{z}})\in\partial_{\rm L}g(\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))$ [23, Lemma 3.2].
@@
    Assume that $g$ is Lipschitz continuous with a Lipschitz constant $L_{g}>0$. By [23, Lemma 3.3] and [69, Prop. 3.4] [^4] In [69, Prop. 3.4], the inequality (23) is proved under the convexity of $g$. However, the proof in [69, Prop. 3.4] can be applied to the weakly convex case as well. , the following hold:
@@
    Moreover, for $0<\mu_{2}<\mu_{1}<\eta^{-1}$, we have [57, Thm. 1.25], [23, Lemma 4.1]
@@
3 Relation between First-order Optimality Conditions of Problems 1.1 and 1.2
This section presents a relation between the original Problem 1.1 and the parameterized Problem 1.2. We begin with the notions of optimality.
@@
    (Stationary point) $\bm{x}^{\star}\in\mathcal{X}$ is said to be a stationary point [^5] A stationary point $\bm{x}^{\star}\in\mathcal{X}$ is also called a limiting stationary point of $J$ in order to distinguish from a Fréchet stationary point where $\partial_{\rm F}J(\bm{x}^{\star})\ni\bm{0}$ holds. However, these stationarities are identical in a case where $J$ is subdifferentially regular in Definition 2.2 (b) (see [57, Cor. 8.11]). Since all functions in this paper are subdifferentially regular, we do not distinguish these stationarities in this paper. of $J$ if $\partial_{\rm L}J(\bm{x}^{\star})\ni\bm{0}$.
By Fermat’s rule [57, Thm. 10.1], for a proper function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, $\bm{x}^{\star\star}\in\mathcal{X}$ is a stationary point of $J$ if $\bm{x}^{\star\star}$ is a local minimizer of $J$. Thus, the condition $\partial_{\rm L}J(\bm{x}^{\star})\ni\bm{0}$ at $\bm{x}^{\star}\in\mathcal{X}$ seems to be the most versatile expressions of the first-order optimality condition for minimization of $J$ at $\bm{x}^{\star}$ (see, e.g., [57, p.207]). For this reason, we focus on finding such stationary points of Problem 1.1 and Problem 1.2 throughout this paper. By Lemma 2.4, the first-order optimality conditions can be expressed respectively with $F=h+g\circ\mathfrak{S}$ as:

### v4

@@ 1 Introduction
    In Section 4.2, we propose a single-loop variable smoothing algorithm for Problem 1.1 via Problem 1.4. For the proposed algorithm, we present an asymptotic convergence guarantee to a stationary point $\bm{y}^{\star}\in\mathcal{Y}$ of Problem 1.4 (Theorem 4.9) by leveraging the gradient consistency property in Theorem 4.4. Moreover, $\bm{\varphi}(\bm{y}^{\star})\in C$ is a stationary point of Problem 1.1 under a certain condition of $\bm{\varphi}$ (see Theorem 3.2 and Corollary 3.4). Even without this condition on $\bm{\varphi}$, $\bm{\varphi}(\bm{y}^{\star})$ satisfies a necessary condition for the local optimality of Problem 1.1 (see Remark 3.6). To the best of the authors’ knowledge, for Problem 1.1 with a nonlinear continuously differentiable mapping $\mathfrak{S}$ and a weakly convex function $g$, the proposed algorithm is the first single-loop algorithm with an asymptotic convergence guarantee to a stationary point of Problem 1.4, and hence to a stationary point of Problem 1.1 under the above condition on $\bm{\varphi}$. In particular, the proposed algorithm reproduces the variable smoothing algorithm [13] in the unconstrained case $C=\mathcal{X}$ (equivalently, $\bm{\varphi}=\mathrm{Id}$ in Problem 1.4) and linear $\mathfrak{S}$, while the proposed algorithm with an asymptotic convergence guarantee covers the case of nonlinear $\mathfrak{S}$ (Note: such an asymptotic convergence guarantee has not been established in [13]; see also Table 1).
  4. (IV)
    As a supplementary result, Corollary 4.12 establishes a first-order oracle complexity $\mathcal{O}(\epsilon^{-3})$ of the proposed algorithm for finding an $\epsilon$-approximate stationary point of Problem 1.4. We note that our approximate stationarity notion in Definition 4.10 (b) is not directly comparable with the notions used in, e.g., [66, 10, 83] and [25, 80], for Problem 1.1 because the underlying stationarity notions differ. Instead, our notion can be viewed as a generalization of the $\epsilon$-approximate stationarity used in the variable smoothing algorithm [13] (see Remark 4.11). More precisely, in a special case $\bm{\varphi}=\mathrm{Id}$ and linear $\mathfrak{S}$, Definition 4.10 (b) reduces to the approximate stationarity notion in [13], and hence our analysis in Corollary 4.12 recovers the oracle complexity $\mathcal{O}(\epsilon^{-3})$ of [13] under that special case.
@@
For a differentiable function $J:\mathcal{X}\to\mathbb{R}$, $\nabla J(\bm{x})\in\mathcal{X}$ is the gradient of $J$ at $\bm{x}\in\mathcal{X}$ if $\mathrm{D}J(\bm{x})[\bm{v}]={\left\langle\nabla J(\bm{x}),\bm{v}\right\rangle}$ for all $\bm{v}\in\mathcal{X}$. Note that for a differentiable mapping $\mathcal{F}:\mathcal{Y}\to\mathcal{X}$, we have the expression [^2] It can be checked by $(\forall\bm{v}\in\mathcal{Y})\ {\left\langle\nabla(J\circ\mathcal{F})(\bm{x}),\bm{v}\right\rangle}_{\mathcal{Y}}=\mathrm{D}(J\circ\mathcal{F})(\bm{x})[\bm{v}]=\mathrm{D}J(\mathcal{F}(\bm{x}))[\mathrm{D}\mathcal{F}(\bm{x})[\bm{v}]]={\left\langle\nabla J(\mathcal{F}(\bm{x})),\mathrm{D}\mathcal{F}(\bm{x})[\bm{v}]\right\rangle}_{\mathcal{X}}={\left\langle(\mathrm{D}\mathcal{F}(\bm{x}))^{*}[\nabla J(\mathcal{F}(\bm{x}))],\bm{v}\right\rangle}_{\mathcal{Y}}$. $\nabla(J\circ\mathcal{F})(\bm{y})=(\mathrm{D}\mathcal{F}(\bm{y}))^{*}[\nabla J(\mathcal{F}(\bm{y}))]\ (\bm{y}\in\mathcal{Y})$. For a function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, $\mathrm{dom}(J)\coloneqq\{\bm{x}\in\mathcal{X}\mid J(\bm{x})<+\infty\}\subset\mathcal{X}$ and $\mathrm{epi}(J):=\{(\bm{x},a)\in\mathcal{X}\times\mathbb{R}\mid J(\bm{x})\leq a\}\subset\mathcal{X}\times\mathbb{R}$ denote respectively the domain and the epigraph of $J$. Moreover, $J$ is said to be (i) proper if $\mathrm{dom}(J)\neq\emptyset$; (ii) lower semicontinuous if $\mathrm{epi}(J)$ is closed in $\mathcal{X}\times\mathbb{R}$; (iii) locally Lipschitz continuous at $\widebar{\bm{x}}\in\mathcal{X}$ if there exist an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ and $L>0$ such that $\lvert J(\bm{x}_{1})-J(\bm{x}_{2})\rvert\leq L\left\lVert\bm{x}_{1}-\bm{x}_{2}\right\rVert\ (\bm{x}_{1},\bm{x}_{2}\in\mathcal{N}_{\widebar{\bm{x}}})$. For a function $J:\mathcal{X}\to\mathbb{R}\cup\{-\infty,+\infty\}$, the limit inferior [69, Definition 1.5] and the limit superior [69, p.13] of $J$ at $\widebar{\bm{x}}\in\mathcal{X}$ are defined respectively by
@@
For a set sequence $(E_{n})_{n=1}^{\infty}$ with subsets $E_{n}\subset\mathcal{X}$, the outer limit [69, Definition 4.1] of $(E_{n})_{n=1}^{\infty}$ is defined by
@@
where $\aleph_{0}$ stands for the cardinality of $\mathbb{N}$. With the outer limit of a set sequence, the outer limit [69, p.152] of a set-valued mapping $S:\mathcal{X}\rightrightarrows\mathcal{X}$ at $\widebar{\bm{x}}\in\mathcal{X}$ is defined as
@@ 2.1 Subdifferential and Normal Cone
where $\partial_{\rm F}J(\widebar{\bm{x}})\coloneqq\left\{\bm{v}\in\mathcal{X}\mid\liminf_{\mathcal{X}\setminus\{\widebar{\bm{x}}\}\ni\bm{x}\to\widebar{\bm{x}}}\frac{J(\bm{x})-J(\widebar{\bm{x}})-{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\geq 0\right\}$ is called the Fréchet (or regular) subdifferential of $J$ at $\widebar{\bm{x}}\in\mathrm{dom}(J)$, and $\partial_{\rm L}J(\widebar{\bm{x}})$ and $\partial_{\rm F}J(\widebar{\bm{x}})$ at $\widebar{\bm{x}}\notin\mathrm{dom}(J)$ are understood as $\emptyset$ (Note: for a convex function $J$, $\partial_{\rm L}J$ and $\partial_{\rm F}J$ are identical to the standard convex subdifferential of $J$ [69, Prop. 8.12]).
By using the subdifferentials, the general normal cone $N_{C}(\widebar{\bm{x}})$ and the regular normal cone $\widehat{N}_{C}(\widebar{\bm{x}})$ to a nonempty subset $C\subset\mathcal{X}$ at $\widebar{\bm{x}}\in C$ are defined (see [69, Def. 6.3, Prop. 6.5, and Ex. 8.14]) respectively by
@@
With notions of normal cone operators, we define regularities of a subset and functions as follows. Such regularities ensure the chain and sum rules for the limiting subdifferential as shown in Lemma 2.4 below (see also Fact B.4 in Appendix B).
@@
    (Clarke regularity [69, Def. 6.4]) A nonempty set $C\subset\mathcal{X}$ is said to be Clarke regular at $\widebar{\bm{x}}\in C$ if (i) $N_{C}(\widebar{\bm{x}})=\widehat{N}_{C}(\widebar{\bm{x}})$ (Note: the inclusion $N_{C}(\widebar{\bm{x}})\supset\widehat{N}_{C}(\widebar{\bm{x}})$ always holds), and (ii) $C\cap\overline{\mathcal{N}}_{\widebar{\bm{x}}}$ is closed for some closed neighborhood [^3] In other words, $\overline{\mathcal{N}}_{\widebar{\bm{x}}}$ is a closed set that contains $\widebar{\bm{x}}$ as an interior point of $\overline{\mathcal{N}}_{\widebar{\bm{x}}}$. $\overline{\mathcal{N}}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$. Moreover, $C$ is said to be Clarke regular if it is Clarke regular at every $\widebar{\bm{x}}\in C$.
@@
    (Subdifferential regularity [69, Def. 7.25]) A proper lower semicontinuous function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ is said to be subdifferentially regular at $\widebar{\bm{x}}\in\mathrm{dom}(J)$ if $\mathrm{epi}(J)\subset\mathcal{X}\times\mathbb{R}$ is Clarke regular at $(\widebar{\bm{x}},J(\widebar{\bm{x}}))$. Moreover, $J$ is said to be subdifferentially regular if it is subdifferentially regular at every $\widebar{\bm{x}}\in\mathrm{dom}(J)$. We note that all functions in this paper are shown to be subdifferentially regular (see, e.g., Lemma 2.4 below and also Example B.2).
A typical Clarke regular set is a closed convex set [69, Thm. 6.9]. Another important example of Clarke regular sets is an embedded submanifold of the Euclidean space $\mathcal{X}$ [69, Exm. 6.8] as follows.
Example 2.3 (Embedded submanifold of $\mathcal{X}$ [69, Exm. 6.8]).
Let $(\emptyset\neq)C\subset\mathcal{X}$ be a $d(\leq\dim(\mathcal{X}))$-dimensional embedded submanifold of $\mathcal{X}$, i.e., for each $\widebar{\bm{x}}\in C$, there exist (i) an open neighborhood $\mathcal{N}_{\widebar{\bm{x}}}\subset\mathcal{X}$ of $\widebar{\bm{x}}$ in $\mathcal{X}$ and (ii) a continuously differentiable mapping $H_{\widebar{\bm{x}}}:\mathcal{N}_{\widebar{\bm{x}}}\to\mathbb{R}^{\dim(\mathcal{X})-d}$ such that $C\cap\mathcal{N}_{\widebar{\bm{x}}}=H_{\widebar{\bm{x}}}^{-1}(\bm{0})$ and $\mathrm{Rank}(\mathrm{D}H_{\widebar{\bm{x}}}(\widebar{\bm{x}}))=\dim(\mathcal{X})-d$, where $H_{\widebar{\bm{x}}}$ is called a local defining map [16, Def. 3.10] (Note: by Whitney’s embedding theorem (see, e.g., [54, Thm. 6.15]), every $d$-dimensional smooth manifold can be seen as an embedded submanifold in $\mathbb{R}^{2d+1}$). Then, by [69, Def. 6.4 and Exm. 6.8], the embedded submanifold $C$ is Clarke regular, and we have the following expression:
@@
where the orthogonal complement subspace $T_{C}(\widebar{\bm{x}})$ of $N_{C}(\widebar{\bm{x}})$ is called the tangent cone to $C$ at $\widebar{\bm{x}}$ (Note: $N_{C}(\widebar{\bm{x}})$ and $T_{C}(\widebar{\bm{x}})$ are also called respectively the normal space and the tangent space to the manifold $C$ at $\widebar{\bm{x}}\in C$ [16, Thm. 3.15, Def. 3.16, and p.107]).
Lemma 2.4 (Subdifferential calculus for Problems 1.1 and 1.4).
Consider $F=h+g\circ\mathfrak{S}$, $C\subset\mathcal{X}$ and $\bm{\varphi}$ in Problems 1.1 and 1.4. Then, $F$, $F+\iota_{C}$ and $F\circ\bm{\varphi}$ are subdifferentially regular, and we have the following expressions:
@@
Fact 2.5 (Normal cone to image sets [69, Thm. 6.43]).
@@
The first inclusion is verified by [69, Thm. 6.43]. The equality follows by the relation $(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\bm{v}]\in\{\bm{0}\}\Leftrightarrow\bm{v}\in\mathrm{Ker}\left((\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}\right)\ (\widebar{\bm{y}}\in\bm{\varphi}^{-1}(\widebar{\bm{x}}))$. ∎
@@ 2.2 Proximity Operator and Moreau Envelope
The proximity operator and the Moreau envelope have been serving as key ingredients for minimization of nonsmooth convex functions (see, e.g., [8, 28]). Even for weakly convex functions, they are useful in design of nonsmooth optimization algorithms. We note that closed-form expressions of the proximity operators for commonly-used weakly convex functions have been found, e.g., $\ell_{1}$-norm [8], Minimax Concave Penalty (MCP) [89], Smoothly Clipped Absolute Deviation (SCAD) [35], log-sum penalty [68], piece-wise exponential function [60], and arctangent penalty [40].
Definition 2.6 (Proximity operator and Moreau envelope [69, Def. 1.22]).
@@
    For any $\widebar{\bm{z}}\in\mathcal{Z}$ and $(\bm{z}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ satisfying $\lim_{n\to\infty}\bm{z}_{n}=\widebar{\bm{z}}$, we have $\lim_{n\to\infty}\mathrm{prox}_{\mu_{n}g}(\bm{z}_{n})=P_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}(\widebar{\bm{z}})$ if $(\mu_{n})_{n=1}^{\infty}\searrow 0$ with $\mu_{n}\in(0,\eta^{-1})$ [7, Lemma 2.3], where $\mathrm{dom}(\partial_{\rm L}g)\coloneqq\{\bm{z}\in\mathcal{Z}\mid\partial_{\rm L}g(\bm{z})\neq\emptyset\}\subset\mathcal{Z}$, $\overline{(\cdot)}$ denotes the closure of a given set, and $P_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}=\mathrm{prox}_{\iota_{\overline{\mathrm{dom}(\partial_{\rm L}g)}}}$ denotes the projection mapping onto the closed convex set $\overline{\mathrm{dom}(\partial_{\rm L}g)}\subset\mathcal{Z}$.
@@
    For any $\mu\in(0,\eta^{-1})$, (i) $^{\mu}g:\mathcal{Z}\to\mathbb{R}$ is continuously differentiable with $\nabla{}^{\mu}g=\frac{\mathrm{Id}-\mathrm{prox}_{\mu g}}{\mu}$, and (ii) $\nabla{}^{\mu}g$ is Lipschitz continuous with a Lipschitz constant $\max\left\{\mu^{-1},\frac{\eta}{1-\eta\mu}\right\}$ [42, Cor. 3.4].
@@
    $(\widebar{\bm{z}}\in\mathcal{Z},\mu\in(0,\eta^{-1}))\quad\nabla{}^{\mu}g(\widebar{\bm{z}})\in\partial_{\rm L}g(\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))$ [13, Lemma 3.2].
@@
    Assume that $g$ is Lipschitz continuous with a Lipschitz constant $L_{g}>0$. By [13, Lemma 3.3] and [67, Prop. 3.4] [^4] In [67, Prop. 3.4], the inequality (23) is proved under the convexity of $g$. However, the proof in [67, Prop. 3.4] can be applied to the weakly convex case as well. , the following hold:
@@
    Moreover, for $0<\mu_{2}<\mu_{1}<\eta^{-1}$, we have [69, Thm. 1.25], [13, Lemma 4.1]
@@
3 Relation between First-order Optimality Conditions of Problems 1.1 and 1.4
This section presents a relation between the original Problem 1.1 and the parameterized Problem 1.4. We begin with the notions of optimality.
@@
    (Stationary point) $\bm{x}^{\star}\in\mathcal{X}$ is said to be a stationary point [^5] A stationary point $\bm{x}^{\star}\in\mathcal{X}$ is also called a limiting stationary point of $J$ in order to distinguish from a Fréchet stationary point where $\partial_{\rm F}J(\bm{x}^{\star})\ni\bm{0}$ holds. However, these stationarities are identical in a case where $J$ is subdifferentially regular in Definition 2.2 (b) (see [69, Cor. 8.11]). Since all functions in this paper are subdifferentially regular, we do not distinguish these stationarities in this paper. of $J$ if $\partial_{\rm L}J(\bm{x}^{\star})\ni\bm{0}$.
By Fermat’s rule [69, Thm. 10.1], for a proper function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$, $\bm{x}^{\star\star}\in\mathcal{X}$ is a stationary point of $J$ if $\bm{x}^{\star\star}$ is a local minimizer of $J$. Thus, the condition $\partial_{\rm L}J(\bm{x}^{\star})\ni\bm{0}$ at $\bm{x}^{\star}\in\mathcal{X}$ seems to be the most versatile expressions of the first-order optimality condition for minimization of $J$ at $\bm{x}^{\star}$ (see, e.g., [69, p.207]). For this reason, we focus on finding such stationary points of Problem 1.1 and Problem 1.4 throughout this paper. By Lemma 2.4, the first-order optimality conditions can be expressed respectively with $F=h+g\circ\mathfrak{S}$ as: